# Anatomical Foundation Models for Brain MRIs

### v3

@@ Abstract
Deep Learning (DL) in neuroimaging has become increasingly relevant for detecting neurological conditions and neurodegenerative disorders. One of the most predominant biomarkers in neuroimaging is represented by brain age, which has been shown to be a good indicator for different conditions, such as Alzheimer’s Disease. Using brain age for weakly supervised pre-training of DL models in transfer learning settings has also recently shown promising results, especially when dealing with data scarcity of different conditions. On the other hand, anatomical information of brain MRIs (e.g. cortical thickness) can provide important information for learning good representations that can be transferred to many downstream tasks. In this work, we propose AnatCL, an anatomical foundation model for brain MRIs that i.) leverages anatomical information in a weakly contrastive learning approach, and ii.) achieves state-of-the-art performances across many different downstream tasks. To validate our approach we consider 12 different downstream tasks for the diagnosis of different conditions such as Alzheimer’s Disease, autism spectrum disorder, and schizophrenia. Furthermore, we also target the prediction of 10 different clinical assessment scores using structural MRI data. Our findings show that incorporating anatomical information during pre-training leads to more robust and generalizable representations. Pre-trained models can be found at: https://github.com/EIDOSLAB/AnatCL.
@@ 1 Introduction
Figure 1: We propose AnatCL, a contrastive learning approach that leverages patient’s metadata (i.e. age) alongside anatomical information derived from structural MRIs, in order to learn meaningful and general representation. Given two input samples $x$ and $x_{i}$, the objective of AnatCL is to learn embeddings $z$ and $z_{i}$, such that their distance in the representation space is proportional to the difference in age and brain anatomy between the two samples.
@@
Contrastive learning methods, particularly those leveraging weakly supervised loss functions, have shown to be effective in creating models that learn robust representation spaces [12, 14]. These methods typically utilize metadata, such as the patient’s age, to guide the learning process, with the aim of aligning similar data points (i.e. similar age) and distinguishing dissimilar ones in the learned representation space. While this approach has yielded promising results, it may fall short of fully capturing the multifaceted information inherent in MRI data and the anatomical complexity of the brain, as it primarily relies on age alone.
@@
For this reason, in this paper, we extend the capabilities of weakly supervised contrastive learning approaches by incorporating additional anatomical measures into the loss formulation. We propose AnatCL, a contrastive learning framework that improves upon previous works leveraging patient age as proxy metadata [12, 1], to also include information derived from anatomical measures (such as cortical thickness), computed on region of interests (ROIs) of common atlas such as the Desikan-Killiany parcellation [7] Specifically, we propose a modified contrastive learning loss function that integrates patient age with a targeted selection of three anatomical features: Mean Cortical Thickness, Gray Matter Volume, Surface Area. This selective integration was designed to assess the effectiveness of using potentially more informative anatomical features in conjunction with demographic data.
We propose and test two versions of our framework: a local version that considers the distance between local anatomical measurements in the Desikan format (a set of $68\times 3$ features), and a global version that considers the same three features across all Desikan regions ($3\times 68$ features). The local version aims to capture the localized anatomical variations between specific ROIs, while the global version aggregates these features across the entire brain to provide a more holistic view of the general anatomy. A graphical overview of AnatCL is provided in Fig. 1.
Figure 2: Overview of the datasets used in this work. We pre-train AnatCL on the large dataset OpenBHB of brain MRIs, and we test different downstream tasks on different conditions such as Alzheimer’s Disease, Schizophrenia, and Autism Spectrum Disorder (ASD).
@@
We pre-train our models using the proposed AnatCL loss on the OpenBHB dataset, a large collection of healthy samples, and then we test the resulting models on different downstream tasks across a variety of publicly available datasets (Sec. 4.1). Our results show that in several downstream tasks, our method can achieve better results than standard state-of-the-art approaches based on brain age prediction. These findings suggest that, in contrastive learning frameworks, incorporating anatomical features alongside the patient’s age can enhance the expressiveness of the learned representations, leading to the development of models that learn more meaningful and generalizable representation spaces. Additionally, we also perform experiments to assess whether anatomical information can be leveraged for predicting phenotypes and clinical assessment scores. To the best of our knowledge, this is one of the first works attempting to tackle this objective. In summary, our contribution is multifold:
@@
    We propose a novel weakly contrastive learning formulation that is able to take into account multiple features and metadata for guiding the learning process;
@@
    We propose AnatCL, a foundation model based on both anatomical and demographic features to build robust and generalizable representations of structural neuroimaging data;
@@
    We perform extensive validation across 12 different downstream clinical tasks (diagnosis) and prediction of 10 different phenotypes and clinical assessment scores.
Furthermore, we also publicly release our pre-trained models for researchers worldwide at https://github.com/EIDOSLAB/AnatCL.
@@ 2 Related Works
Machine learning on structural brain imaging. Current literature has been mainly devoted to segmentation tasks for brain anatomical data, such as brain tumor segmentation [2] or tissue segmentation [10], in a supervised setting requiring annotated maps. The deep neural network U-Net [27] and its extension nnU-Net [19] were very successful at such tasks. Another body of literature tackled phenotype and clinical prediction tasks from structural imaging, such as age prediction [15], Alzheimer’s disease detection [33], cognitive impairment detection, or psychiatric conditions prediction [24, 4]. Current approaches mostly involve traditional machine learning models such as SVM or penalized linear models trained in a supervised fashion to predict the targeted phenotypes. Since most studies only include a few hundred subjects in their datasets, larger models introducing more parameters (such as neural networks) do not necessarily translate into better performance [14, 28, 17], probably because of over-fitting.
Self-supervised learning on brain imaging. Self-supervised learning has been popularized by its impressive performance in Natural Language Processing (NLP) using auto-regressive models trained on web-scale text corpus and mainly deployed on generation tasks [8, 3]. Since it does not require a supervised signal for training, it is particularly appealing when large-scale data are available, and the annotation cost is prohibitive. Its applicability on neuroimage data is still unclear and very few studies have tackled this problem so far [18].  [12] demonstrated that contrastive learning integrating subjects meta-data can improve the classification of patients with psychiatric conditions.  [31] demonstrated that learning frameworks from NLP, such as sequence-to-sequence autoencoding, causal language modeling, and masked language modeling, can improve mental state decoding from brain activity recorded with functional MRI.
Contrastive Learning on brain imaging Recently, interests in contrastive learning approaches for brain age prediction has risen. It has been demonstrated that contrastive learning performs better that traditional methods for neuroimaging [14]. Among different works we can identify y-Aware [12] and ExpW [1]. These works propose to utilize contrastive learning for continuous regression tasks, by employing a kernel to measure similarity between different samples. In these works, age is used as pre-training target, followed by transfer learning on specific diseases. However, they are limited to only considering age as pretraining feature. In this work, we extend this formulation to work with multiple attributes.
@@ 3.1 Background
Contrastive learning (CL) (either self-supervised as in SimCLR [5] or supervised as in SupCon [21]) leverages discrete labels (i.e., categories) to define positive and negative samples. Starting from a sample $x$, called the anchor, and its latent representation $z=f(x)$, CL approaches aim at aligning the representations of all positive samples to $z$, while repelling the representations of the negative ones. CL is thus not adapted for regression (i.e., continuous labels), as a hard boundary between positive and negative samples cannot be determined. Recently, there have been developments in the field of CL for tackling regression problems, especially aimed at brain age prediction, such as y-Aware [12] and ExpW [1], from brain MRIs. These approaches can be broadly categorized as weakly contrastive, as brain age is used as pre-train task.
Figure 3: Graphical evaluation of AnatCL (local and global) on all tasks considered. Overall, AnatCL achieves higher results more consistently than any other baselines, with notable performance on ASD and Schizophrenia.
3.2 Weakly contrastive learning
To tackle the issue of continuous labels, weakly contrastive approaches employ the notion of a degree of “positiveness” between samples. This degree is usually defined by a kernel function $w_{i}=\mathcal{K}(y-y_{i})$, where $0\leq w_{i}\leq 1$ is computed by a Gaussian or a Radial Basis Function (RBF) kernel and $y,y_{i}$ are the continuous attributes of interest (e.g. age). The goal of weakly contrastive learning is thus to learn a parametric function $f:\mathcal{X}\rightarrow\mathbb{S}^{d}$ that maps samples with a high degree of positiveness ($w_{i}\sim 1$) close in the latent space and samples with a low degree ($w_{i}\sim 0$) far away from each other. y-Aware [12] proposes to do so by aligning positive samples with a strength proportional to the degree, i.e.:
@@ 3.2 Weakly contrastive learning
where $z=f(x)$, $z_{i}=f(x_{i})$, $z_{t}=f(x_{t})$, and $sim$ is a similarity function (e.g. cosine). A similar approach is also adopted by ExpW [1]. A limitation of this class of approaches is the reliance on one single attribute (age) to determine the alignment strength. They are thus not suited to leverage multiple attributes (e.g. anatomical measurements of the brain). Our proposed approach aims to solve this issue by extending the weakly contrastive formulation to include multiple attributes.
@@ 3.3 Anatomical Contrastive Learning
We propose a novel formulation for weakly contrastive learning to employ multiple anatomical attributes derived from the MRI scan, that we call AnatCL. The features we consider are derived from the standard Desikan-Kiliany parcellation [7], based on the information available in the OpenBHB dataset (see Sec. 4.1), composed of $K=68$ ROIs. Among the anatomical measurements, we consider cortical thickness (CT), gray matter volume (GMV), and surface area (SA) [13]. To include these measurements inside AnatCL, we propose two different formulation, global and local, based on how the similarity across two brain scans is computed.
3.3.1 Local Descriptor
Given a sample $x$, we consider the set of local descriptors $\Psi(x)=\{\psi(x)^{1},\dots\,\psi(x)^{K}\}$, where $\psi(x)^{k}\in\mathbb{R}^{3}$ for $k\in\{1,\dots,K\}$, and $K$ is the number of region of interest (ROIs) in the chosen atlas (Desikan). Each local descriptor is thus a vector composed of the three anatomical measurements CT, GMV, and SA, corresponding to a specific ROI of the atlas. In order to compute a degree of positiveness $\alpha_{i}$ based on the local descriptors of two samples $x$ and $x_{i}$, we consider the average cross-region similarity as:
@@ 3.3.1 Local Descriptor
where, for brevity $\psi^{k}=\psi(x)^{k}$ and $\psi_{i}^{k}=\psi(x_{i})^{k}$, $s$ can either be a similarity function or a kernel as in [12], and $\gamma(\psi)$ is a normalization function that normalizes each component of $\psi$ to a standard range $[0,1]$. In this work, for simplicity, we choose to employ cosine similarity as $s$. By plugging $a_{i}$ into Eq. 1, we obtain the local variant of AnatCL:
@@
where we omitted the normalization terms $1/\sum_{j}\alpha_{k}$ and $1/K$ for brevity.
3.3.2 Global Descriptor
Taking a complementary approach, we now consider global anatomy descriptors $\omega(x)^{j}\in\mathbb{R}^{K}$ for the entire brain with $j\in\{1,2,3\}$, that contain the values across all $K=68$ regions for each anatomical measurement. Similarly to above, in order to compute a degree of positiveness $\beta_{i}$ between two samples $x$ and $x_{i}$ given their global anatomical descriptors $\Omega(x)=\{\omega(x)^{1},\dots,\omega(x)^{3}\}$ and $\Omega(x_{i})=\{\omega(x_{i})^{1},\dots,\omega(x_{i})^{3}\}$ we compute the cross-measurement similarity (global descriptor) as:
$\beta_{i}=\frac{1}{3}\sum_{j=1}^{3}s(\omega^{j},\omega_{i}^{j})$ (4)
where, similarly to before, $\omega^{j}=\omega(x)^{j}$ and $\omega_{i}^{j}=\omega(x_{i})^{j}$. This time we do not need to normalize $\omega^{j}$ and $\omega_{i}^{j}$ as $s$ is evaluated between features of the same type, so the scale of the values is comparable. Also in this case, for the sake of simplicity, we employ a cosine similarity function for $s$. As before, by plugging $\beta_{i}$ in Eq. 1 we obtain the AnatCL global loss:
$\begin{split}\mathcal{L}_{AnatCL-G3}=\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\\ =-\sum_{i}\sum_{j=1}^{3}s(\omega^{j},\omega_{i}^{j})\log\left(\frac{\exp(sim(z,z_{i}))}{\sum_{t}^{N}\exp(sim(z,z_{t}))}\right)\end{split}$ (5)
where, as before, normalization terms are omitted.

### v4

@@ Abstract
Deep Learning (DL) in neuroimaging has become increasingly relevant for detecting neurological conditions and neurodegenerative disorders. One of the most predominant biomarkers in neuroimaging is represented by brain age, which has been shown to be a good indicator for different conditions, such as Alzheimer’s Disease. Using brain age for weakly supervised pre-training of DL models in transfer learning settings has also recently shown promising results, especially when dealing with data scarcity of different conditions. On the other hand, anatomical information of brain MRIs (e.g. cortical thickness) can provide important information for learning good representations that can be transferred to many downstream tasks. In this work, we propose AnatCL, an anatomical foundation model for structural brain MRIs that i.) leverages anatomical information in a weakly contrastive learning approach, and ii.) achieves state-of-the-art performances across many different downstream tasks. To validate our approach we consider 12 different downstream tasks for the diagnosis of different conditions such as Alzheimer’s Disease, autism spectrum disorder, and schizophrenia. Furthermore, we also target the prediction of 10 different clinical assessment scores using structural MRI data. Our findings show that incorporating anatomical information during pre-training leads to more robust and generalizable representations. Pre-trained models can be found at: https://github.com/EIDOSLAB/AnatCL.
keywords:
Brain foundation models , Contrastive Learning , Neuroimaging
journal: Pattern Recognition Letters [^1] Data used in preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). As such, the investigators within the ADNI contributed to the design and implementation of ADNI and/or provided data but did not participate in the analysis or writing of this report. A complete listing of ADNI investigators can be found at: http://adni.loni.usc.edu/wp-content/uploads/how_to_apply/ADNI_Acknowledgement_List.pdf
[…]
[unito] organization=University of Turin,city=Turin, country=Italy
[…]
[neurospin] organization=GAIA, BAOAB, NeuroSpin, CEA Saclay, city=Saclay, country=France
@@ 1 Introduction
Figure 1: We propose AnatCL, a contrastive learning approach that leverages patient’s metadata (i.e., age) alongside anatomical information derived from structural MRIs, in order to learn meaningful and general representations. Given two input samples $x$ and $x_{i}$, the objective of AnatCL is to learn embeddings $z$ and $z_{i}$, such that their distance in the representation space is proportional to the difference in age and brain anatomy between the two samples.
@@
Contrastive learning methods, particularly those leveraging weakly supervised loss functions, have been shown to be effective in creating models that learn robust representation spaces (Dufumier et al., 2021, 2024). These methods typically utilize metadata, such as the patient’s age, to guide the learning process, to align similar data points (e.g., similar age) and distinguish dissimilar ones in the learned representation space. While this approach has yielded promising results, it may fall short of fully capturing the multifaceted information inherent in MRI data and the anatomical complexity of the brain, as it primarily relies on age alone.
@@
In this paper, we extend the capabilities of weakly supervised contrastive learning approaches by incorporating additional anatomical measures into the loss formulation. We propose AnatCL, a contrastive learning framework that improves upon previous works leveraging patient age as proxy metadata (Dufumier et al., 2021; Barbano et al., 2023), to also include information derived from anatomical measures (such as cortical thickness), computed on region of interests (ROIs) of common atlas such as Desikan-Killiany (Desikan et al., 2006) or Destrieux (Destrieux et al., 2010). Specifically, inspired by the MIND anatomical similarity metric (Sebenius et al., 2023), we propose a modified contrastive learning loss function that integrates patient age with a targeted selection of three anatomical features: Mean Cortical Thickness, Gray Matter Volume, Surface Area. This selective integration was designed to assess the effectiveness of using potentially more informative anatomical features in conjunction with demographic data.
@@
We propose and test two versions of our framework: a local version and a global version. The local version aims to capture the localized anatomical variations between specific ROIs (i.e., computing a cross-region similarity across all measures), while the global version aggregates these features across the entire brain to provide a more holistic view of the general anatomy (i.e., computing a cross-measure similarity across all regions). A graphical overview of AnatCL is provided in Fig. 1. We pre-train our models using the proposed AnatCL loss on the OpenBHB dataset, a large collection of healthy samples, and then we test the resulting models on different downstream tasks across a variety of publicly available datasets (Sec. 4.1). Our results show that in several downstream tasks, AnatCL achieves better results than traditional pre-training approaches based on brain age prediction only. These findings suggest that, in contrastive learning frameworks, incorporating anatomical features alongside the patient’s age can enhance the expressiveness of the learned representations, leading to the development of models that learn more meaningful and generalizable representation spaces. Additionally, we also perform experiments to assess whether the embeddings learned using anatomical information can be leveraged for predicting clinical assessment scores, cognition, and psychopathology of the subjects (e.g., IQ, depression scores, and abnormal involuntary movements). To the best of our knowledge, this is one of the first works attempting to tackle this objective, which is typically performed on functional MRI data (Kong et al., 2023). In summary:
@@
    We propose a novel weakly contrastive learning formulation that takes into account multiple features and metadata for guiding the learning process;
@@
    We propose AnatCL, a foundation model based on both anatomical and age features to build robust and generalizable representations of structural neuroimaging data;
@@
    We perform extensive validation across 12 different deep phenotyping tasks and the prediction of 10 different clinical assessment scores.
@@ 2 Related Works
Machine learning on structural brain imaging
Current literature has been mainly devoted to segmentation tasks for brain anatomical data, such as brain tumor segmentation (Biratu et al., 2021) or tissue segmentation (Dora et al., 2017), in a supervised setting requiring annotated maps. The deep neural network U-Net (Ronneberger et al., 2015) and its extension nnU-Net (Isensee et al., 2021) were very successful at such tasks. Another body of literature tackled phenotype and clinical prediction tasks from structural imaging, such as age prediction (Franke & Gaser, 2019), Alzheimer’s disease detection (Wen et al., 2020), cognitive impairment detection, or psychiatric conditions prediction (Nunes et al., 2020; Chand et al., 2020). Current approaches mostly involve traditional machine learning models such as SVM or penalized linear models trained in a supervised fashion to predict the targeted phenotypes. Since most studies only include a few hundred subjects in their datasets, larger models introducing more parameters (such as neural networks) do not necessarily translate into better performance (Dufumier et al., 2024; Schulz et al., 2020; He et al., 2020), probably because of over-fitting.
Self-supervised learning on brain imaging
Self-supervised learning has been popularized by its impressive performance in Natural Language Processing (NLP) using auto-regressive models trained on web-scale text corpus and mainly deployed on generation tasks (Devlin et al., 2018; Brown et al., 2020). Since it does not require a supervised signal for training, it is particularly appealing when large-scale data are available, and the annotation cost is prohibitive. Its applicability on neuroimage data is still unclear and very few studies have tackled this problem so far (Huang et al., 2023). Dufumier et al. (2021) demonstrated that contrastive learning integrating subjects meta-data can improve the classification of patients with psychiatric conditions. Thomas et al. (2022) demonstrated that learning frameworks from NLP, such as sequence-to-sequence autoencoding, causal language modeling, and masked language modeling, can improve mental state decoding from brain activity recorded with functional MRI.
Contrastive Learning on brain imaging
Recently, interest in contrastive learning approaches for brain age prediction has risen. It has been demonstrated that contrastive learning performs better than traditional methods for neuroimaging (Dufumier et al., 2024). Among different works, we can identify y-Aware (Dufumier et al., 2021) and ExpW (Barbano et al., 2023). These works propose to utilize contrastive learning for continuous regression tasks by employing a kernel to measure similarity between different samples. In these works, age prediction is used as pre-training task, followed by transfer learning on specific diseases. However, they are limited to only considering age as pre-training feature. In this work, we extend this formulation to work with multiple attributes.
@@ 3.1 Background
Contrastive learning (CL), either self-supervised as in SimCLR (Chen et al., 2020) or supervised as in SupCon (Khosla et al., 2020), leverages discrete labels (i.e., categories) to define positive and negative samples. Starting from a sample $x$, called the anchor, and its latent representation $z=f(x)$, CL approaches aim at aligning the representations of all positive samples to $z$, while repelling the representations of the negative ones. CL is thus not adapted for regression (i.e., continuous labels), as a hard boundary between positive and negative samples cannot be determined. Recently, there have been developments in the field of CL for tackling regression problems, especially aimed at brain age prediction, such as y-Aware (Dufumier et al., 2021) and ExpW (Barbano et al., 2023), from brain MRIs. These approaches can be broadly categorized as weakly contrastive, as brain age is used as pre-train task.
Figure 2: Graphical evaluation of AnatCL (local and global) on all deep phenotyping tasks considered. Overall, AnatCL achieves higher results more consistently than any other baseline, with notable performance on ASD and Schizophrenia.
3.2 Weakly contrastive learning with age
To tackle the issue of continuous labels, weakly contrastive approaches employ the notion of a degree of “positiveness” between samples (Barbano et al., 2023). This degree is usually defined by a kernel function $w_{i}=\mathcal{K}(y-y_{i})$, where $0\leq w_{i}\leq 1$ is computed by a Gaussian or a Radial Basis Function (RBF) kernel and $y,y_{i}$ are the continuous attributes of interest (e.g. age). The goal of weakly contrastive learning is thus to learn a parametric function $f:\mathcal{X}\rightarrow\mathbb{S}^{d}$ that maps samples with a high degree of positiveness ($w_{i}\sim 1$) close in the latent space and samples with a low degree ($w_{i}\sim 0$) far away from each other. y-Aware (Dufumier et al., 2021) proposes to do so by aligning positive samples with a strength proportional to the degree, i.e.:
@@ 3.2 Weakly contrastive learning with age
where $z=f(x)$, $z_{i}=f(x_{i})$, $z_{t}=f(x_{t})$, and $sim$ is a similarity function (e.g. cosine). A similar approach is also adopted by ExpW (Barbano et al., 2023). A limitation of these approaches is the reliance on one single attribute (age) to determine the alignment strength. They are thus not suited to leverage multiple attributes (e.g., anatomical measurements of the brain). Our proposed approach aims to solve this issue by extending the weakly contrastive formulation to include multiple attributes.
@@ 3.3 Anatomical Contrastive Learning
We propose a novel formulation for weakly contrastive learning to employ anatomical attributes derived from the MRI scan, which we call AnatCL. The features we consider are available in the OpenBHB dataset (see Sec. 4.1) for both Desikan-Killany and Destrieux atlases, composed of 68 and 148 ROIs respectively. Inspired by MIND (Sebenius et al., 2023), we consider cortical thickness (CT), gray matter volume (GMV), and surface area (SA). Additional measures are available in the dataset; however, we did not achieve better results when using them (see supp. material). To include these measurements inside AnatCL, we propose two different formulations, global and local, based on how the similarity across two brain scans is computed.
Local Descriptor
Given a sample $x$, we consider the set of local descriptors $\Psi(x)=\{\psi^{1}(x),\dots\,\psi^{K}(x)\}$, where $\psi^{k}(x)\in\mathbb{R}^{N}$ for $k\in\{1,\dots,K\}$, $K$ is the number of region of interest (ROIs) in the chosen atlas (68 for Desikan, 148 for Destrieux), and $N$ is the number of features (i.e., 3). Each local descriptor is thus a vector composed of the three anatomical measurements CT, GMV, and SA, corresponding to a specific ROI of the atlas. To compute a degree of positiveness $\alpha_{i}$ based on the local descriptors of two samples $x$ and $x_{i}$, we consider the average cross-region similarity as:
@@
where, for brevity $\psi^{k}=\psi^{k}(x)$ and $\psi_{i}^{k}=\psi^{k}(x_{i})$, $s$ can either be a similarity function or a kernel as in Dufumier et al. (2021), and $\gamma(\psi)$ is a normalization function that normalizes each component of $\psi$ to a standard range $[0,1]$. In this work, for simplicity, we choose to employ cosine similarity as $s$. By plugging $a_{i}$ into Eq. 1, we obtain the local variant of AnatCL (omitted the normalization terms $1/\sum_{j}\alpha_{k}$ and $1/K$ for brevity):
@@
Global Descriptor
Taking a complementary approach, we now consider global anatomy descriptors $\omega^{j}(x)\in\mathbb{R}^{K}$ for the entire brain with $j\in\{CT,GMV,SA\}$, that contain the values across all $K$ regions for each anatomical measurement. Similarly to above, in order to compute a degree of positiveness $\beta_{i}$ between two samples $x$ and $x_{i}$ given their global anatomical descriptors $\Omega(x)=\{\omega^{CT}(x),\omega^{GMV}(x),\omega^{SA}(x)\}$ and $\Omega(x_{i})=\{\omega^{CT}(x_{i}),\omega^{GMV}(x_{i}),\omega^{SA}(x_{i})\}$ we compute the cross-measurement similarity (global descriptor) as:
$\beta_{i}=\frac{1}{3}\sum_{j}s(\omega^{j},\omega_{i}^{j})\quad j\in\{CT,GMV,SA\}$ (4)
where, similarly to before, $\omega^{j}=\omega^{j}(x)$ and $\omega_{i}^{j}=\omega^{j}(x_{i})$. This time we do not need to normalize $\omega^{j}$ and $\omega_{i}^{j}$ as $s$ is evaluated between features of the same type, so the scale of the values is comparable. Also in this case, we employ a cosine similarity function for $s$. As before, by plugging $\beta_{i}$ in Eq. 1 we obtain the AnatCL global loss (omitting the normalization terms):
$\begin{split}\mathcal{L}_{AnatCL-G3}=\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\\ =-\sum_{i}\sum_{j}s(\omega^{j},\omega_{i}^{j})\log\left(\frac{\exp(sim(z,z_{i}))}{\sum_{t}^{N}\exp(sim(z,z_{t}))}\right)\end{split}$ (5)